# Expansion Span: Combining Fading Memory and Retrieval in Hybrid State Space Models

### v1

@@ 4.2 Evaluating with Full Attention
Similar to [3], we utilize SE-Attn for efficient fine-tuning and revert to using Full-Attn during evaluation. Due to KV-caching, the complexity of Full-Attn scales linearly during evaluation, so an efficient Attention layer in this setting is not as crucial as in training. Nevertheless, for the sake of completeness, we begin by evaluating our models with the same Attention layer used during fine-tuning (i.e., we fine-tune and deploy our Hybrid model with SE-Attn). In Tab. 1, we experiment with models that use Full-Attn, SW-Attn, $S^{2}$-Attn, and SE-Attn; then, we evaluate perplexity on the PG-19 dataset. All models were fine-tuned with a context size of 8192. We observe that SW-Attn is best at preserving the perplexity on context sizes up to 32×32$\times$32 × larger than the one used for pre-training. SE-Attn is also able to maintain a lower perplexity across longer context sizes, while $S^{2}$-Attn deteriorates much more quickly.
@@
Table 1: Mamba-2-Hybrid fine-tuned with SE-Attn or SW-Attn preserves perplexity up to 32×32$\times$32 × the pre-training context size. We fine-tune a Mamba-2-Hyrbid model (pre-trained on a context size of 2048) on a context size of 8192 using various Attention variants. Then, we deploy them on longer context sizes using the same Attention variant used during adaptation.
4.3 Mamba-2-Hybrid
All of our Mamba-2-Hybrid models are fine-tuned with a context size of 8192. For SE-Attn, we use a block size of 32, and a top-$k$ of 8. For SW-Attn, we use a window size of 4096. $S^{2}$-Attn uses the parameters from [3]. When fine-tuning Mamba-2-Hybrid using SE-Attn, we found that applying the same chunk sizes at each layer leads to suboptimal downstream performance, we therefore segment each sample into chunks of variable sizes (picked randomly from $\{2048,4096\}$). We found this prevents the model from ‘overfitting’ the given segmentation of the past which is an hyper-parameter of our method fixed a-priori and not dependent on the actual data content. An ablation on SE-Attn with different chunk sizes is provided in Sec. 4.5.
Perplexity. We provide PG-19 PPL results for Mamba-2-Hybrid in Tab. 2. All fine-tuned models improve upon the non-fine-tuned model. SE-Attn yields the closest performance to the paragon model fine-tuned with Full-Attn, outperforming $S^{2}$-Attn and SW-Attn across all context sizes at or above the fine-tuning size.
LM Harness. Next, we evaluate Mamba-2-Hybrid models across short and long-context tasks in the LM Evaluation Harness suite. Our results are provided in Tab. 2, where we observe that all models perform similarly on short-context tasks—including the non-fine-tuned model—suggesting there is no performance regression when fine-tuning on larger contexts. Furthermore, we observe that fine-tuning with SE-Attn gives the closest performance to fine-tuning with Full-Attn.
Table 2: Fine-tuning Mamba-2-Hybrid with SE-Attn outperforms fine-tuning with $S^{2}$-Attn and SW-Attn on long-context natural language tasks. We fine-tune Mamba-2-Hybrid with a context size of 8192 using various Attention variants. We evaluate PG-19 validation perplexity (PPL) and observe that fine-tuning with SE-Attn preserves performance at longer contexts better than $S^{2}$-Attn and SW-Attn. On short-context tasks from the LM Harness suite, all models perform similarly. On long context tasks from the LM Harness suite, SE-Attn outperforms $S^{2}$-Attn and SW-Attn. Abbreviations: Hella=HellaSwag, LAMB=LAMBADA, WG=WinoGrande, SQA=ScrollsQAsper, SNQA=ScorllsNarrativeQA.
RULER. Next, we assess the performance of our models on long-context tasks using the RULER benchmark. We plot the average accuracy across eleven RULER tasks in Figure 2(a) (see Sec. 11.4 for details on how we aggregate these metrics). We observe that fine-tuning with SE-Attn yields a performance similar to fine-tuning with Full-Attn, and outperforms $S^{2}$-Attn and SW-Attn. We provide additional RULER metrics in Figure 8, where we see a substantial improvement on the variable tracking (VT) task, which may be attributed to SE-Attn’s retrieval during fine-tuning.
(a) Mamba-2-Hybrid
(b) Llama1
Figure 2: Fine-tuning with SE-Attn outperforms SW-Attn and $S^{2}$-Attn on the RULER benchmark when applied to Mamba-2-Hybrid and Llama1. We fine-tune Mamba-2-Hybrid with a context size of 8192, and Llama1 with 16384 using various Attention variants. We average over eleven RULER tasks, as explained in Sec. 11.4. Fine-tuning with SE-Attn consistently outperforms SW-Attn and $S^{2}$-Attn even when evaluating on context sizes beyond the fine-tuning size.
4.4 Llama1 7B
When fine-tuning Llama 7B models, we found that using a fixed chunk size of $M=4096$ gave better downstream performance. All of our Llama models are fine-tuned with a context size of 16384. For SE-Attn, we use block size of 32, and a top-$k$ of 8. For SW-Attn, we use a window size of 4096, and $S^{2}$-Attn uses the default parameters in [3]. All Attention variants use the same RoPE [29] scaling as in [3].
Perplexity. Fine-tuning Llama with SE-Attn leads to better generalization on context sizes smaller and larger than the one used for fine-tuning. In Tab. 3, we observe that fine-tuning with SE-Attn preserves the performance of the non-fine-tuned model at smaller context sizes, and offers greater generalization to larger context sizes than $S^{2}$-Attn and SW-Attn, as measured on PG-19 validation perplexity.
LM Harness. Fine-tuning Llama with SE-Attn yields a stronger performance on long-context tasks on the LM Evaluation Harness benchmark. As shown in Tab. 3, fine-tuning with SW-Attn, $S^{2}$-Attn, and SE-Attn all improve upon the non-fine-tuned model on shorter context tasks and perform similarly. However, SE-Attn gives a greater performance on long context tasks.
RULER. Fine-tuning Llama with SE-Attn produces a model with stronger performance on RULER tasks than fine-tuning with SW-Attn and $S^{2}$-Attn, as illustrated in Figure 2(b). On average, SW-Attn and $S^{2}$-Attn perform similarly, however, fine-tuning with SE-Attn improves performance by $\sim$5%. Interestingly, despite all models having a similar PPL up to a context size of 16k as shown in Tab. 3, we observe a substantial difference between the RULER performance of models fine-tuned with our SE-Attn and those fine-tuned with SW-Attn and $S^{2}$-Attn . We expand on this observation for Hybrid models in Sec. 6.
4.5 Ablations on Mamba-2-Hybrid
In this section, we ablate over some of the design choices of SE-Attn on Mamba-2-Hybrid. For this analysis, we consider the RULER benchmark, as it is a strong indicator of performance on long-context tasks. As our metric, we use the average of the eleven RULER tasks defined in Sec. 11.4.
Does retrieval during training help? In Sec. 3.2, we introduced SE-Attn-NoMem, a variant of SE-Attn where we do not do any retrieval and process chunks independently. Naturally, we do not expect this to do well due to the lack of shared information across chunks. We confirm this in Figure 3(a), where we observe that SE-Attn-NoMem achieves a much lower performance than SE-Attn with retrieval. Furthermore, we also fine-tune with SE-Attn-Random, a variant of SE-Attn where we retrieve random memory blocks from the past. We observe that this improves upon SE-Attn-NoMem, indicating that retrieving some information from the past improves upon no information. However, SE-Attn-Random does not do as well as SE-Attn, which decides which blocks to retrieve based on relevancy.
Chunk size. For Mamba-2-Hybrid, we found that using a random chunk size during each forward pass improved upon having a fixed chunk size. In Figure 3(b), we compare the performance of SE-Attn, which chooses a random chunk size in $\{2048,4096\}$ during each forward call, to SE-Attn (2048M) and SE-Attn (4096M), which use a fixed chunk size of 2048 and 4096, respectively. We observe that SE-Attn with a fixed chunk size of 4096 improves upon SE-Attn with a fixed chunk size of 2048, but SE-Attn outperforms both. This suggests that training with random chunk sizes may have a regularizing effect, making the model more robust to changes in the input context size.
Figure 3: SE-Attn ablations on Mamba-2-Hybrid. We plot the average of eleven RULER tasks as described in Sec. 11.4. (a): SE-Attn-NoMem processes chunks of tokens without retrieval, while SE-Attn-Random populates its expansion span by retrieving random memory blocks for each chunk. We observe that our Attention-based memory retrieval (SE-Attn) gives the strongest performance. (b): Using SE-Attn with a chunk size chosen randomly from $\{2048,4096\}$ acts as a regularizer and outperforms SE-Attn with fixed chunk sizes of 2048 and 4096. (c): SE-Attn with larger memory blocks (i.e., more tokens per memory block) with a smaller top-$k$ tends to do better than smaller block sizes with a larger top-$k$. (d): An expansion span consisting of 256 total tokens (8 memory blocks with 32 tokens in each) gives the strongest performance. 32S/32k is omitted due to memory constraints.
Block size and top-$k$. For a fixed expansion span size, we might expect that retrieving memory blocks at a finer granularity should perform better than retrieving a smaller number of large blocks; this is because retrieving a larger number of small blocks is as flexible as retrieving a small number of large blocks. However, as illustrated in Figure 3(c), we found that fine-tuning SE-Attn with larger block sizes and a smaller top-k gave the best performance. This is possibly due to the increased complexity of the retrieval task, which gets more difficult as the number of possible blocks to retrieve increases. As we are fine-tuning with LoRA, we hypothesize that the model’s capacity may not be sufficient to learn to retrieve effectively in this setting. As illustrated in Figure 3(d), we found that an expansion span populated with 256 retrieved tokens (32 memory blocks with 8 tokens in each) works best.
4.6 HyLoRA for Hybrid Models
We fine-tune our hybrid models using HyLoRA, which builds upon LoRA+ [3] by also training 1D convolution layers. As illustrated in Figure 4, when evaluated on long-context RULER tasks, fine-tuning SE-Attn models with HyLoRA gives the strongest performance. See Sec. 6 for additional results when fine-tuning with Full-Attn, and the effect of the LoRA rank on downstream performance.
Figure 4: HyLoRA outperforms LoRA and LoRA+ on Hybrid models. We fine-tune Mamba-2-Hybrid with SE-Attn using LoRA, LoRA+, and HyLoRA. We find that LoRA and LoRA+ perform sub-optimally. HyLoRA augments LoRA+ by additionally training the 1D convolution layers and yields strong performance.
4.7 Empirical Runtime Analysis
SE-Attn is much faster than Full-Attn and $S^{2}$-Attn on large contexts. In Figure 5, we profile different Attention layers on various context sizes and see that SE-Attn is much faster than $S^{2}$-Attn and Full-Attn. Moreover, we see that the runtime of SE-Attn is similar to that of SW-Attn, indicating that our retrieval overhead is minimal despite the much longer Attention span. See Sec. 10 for a more detailed runtime analysis of SE-Attn.
Figure 5: SE-Attn is faster than $S^{2}$-Attn and Full-Attn, especially on long contexts. The retrieval overhead of SE-Attn is minimal, with runtime similar to SW-Attn (which has a limited pre-determined Attention span).
5 Conclusion
We close with the limitations of our method. Although we have conducted experiments at a relatively large model scale (2.7B for Hybrid models and 7B for Transformers) and long contexts (up to 32k), testing our method for efficiently adapting larger models on longer contexts is paramount. Furthermore, while we utilize datasets that require modeling of long-range dependencies, we found that perplexity-based tasks do not faithfully measure models’ capabilities to handle long contexts, and instead tasks like RULER provide better signals of long-context capabilities. However, RULER is mostly a synthetic dataset and does not cover more nuanced tasks that require reasoning over long documents. Validating our method on more complex long-range benchmarks is a promising area for future work.
Supplementary Material
6 LoRA for Hybrid Models and the Pitfalls of Perplexity
@@
In this section, we conduct a more thorough ablation study on the effect of training Mamba-2-Hybrid with different variants of LoRA.
@@
Fine-tuning with Full-Attn. We first expand upon Figure 4, where we showed that fine-tuning Mamba-2-Hybrid with SE-Attn using HyLoRA produced a stronger model than training with LoRA and LoRA+. In Figure 6, we provide a similar plot, but using Full-Attn during fine-tuning rather than SE-Attn. We observe a similar trend, where HyLoRA produces a stronger model than LoRA and LoRA+. In the remaining analyses in this section, we provide results for fine-tuning with SE-Attn, as we observed similar trends when fine-tuning with Full-Attn.
@@
Effect of LoRA rank. We next explore how the rank used for LoRA fine-tuning affects downstream performance on the RULER task. Given its stronger performance, we focus on our HyLoRA. In our previous Mamba-2-Hybrid experiments, we used a LoRA rank $(r)$ of 32 and an $\alpha$ of 64 (for our Transformer experiments, we used $r=8$ and $\alpha=16$ as in [3]). In Figure 7, we plot the average RULER results for Mamba-2-Hybrid models fine-tuned with HyLoRA using different ranks (we scale $\alpha$ to maintain the ratio $\alpha/r=2$). Here, we observe that training with a larger rank improves downstream performance, and we start to see some saturation around $r=64$.
In Tab. 4, we provide PG-19 perplexity results for Mamba-2-Hybrid trained with different LoRA variants. All models are fine-tuned with a context size of 8192, and evaluated with multiple context sizes. For a given context size, we do not see a substantial difference in the perplexity. This is similar to the observation in [3]. However, as discussed above, the rank can have a significant effect on downstream long-context tasks that require strong recall capabilities, as in RULER. Hence, while perplexity results may be promising, they are not necessarily indicative of performance on more complex long-context tasks.
@@
Figure 7: Fine-tuning with a larger LoRA rank using HyLoRA improves performance on the RULER benchmark. We fine-tune Mamba-2-Hybrid using HyLoRA with different LoRA ranks (we maintain a LoRA rank to alpha ratio of 2). We observe that fine-tuning with a larger rank produces stronger downstream results on RULER, with some saturation with a rank of 64.
@@
Table 4: Mamba-2-Hybrid LoRA Ablations. We fine-tune Mamba-2-Hybrid with a context size of 8192 with SE-Attn using different LoRA variants. We consider LoRA, LoRA+, and HyLoRA (ours) and evaluate perplexity on the PG-19 dataset. We observe that all LoRA variants yield similar perplexity results. However, as depicted in Figure 4 and Figure 7, different LoRA variants yield substantially different performances on more complex tasks.
7 Additional Fine-Tuning Results
@@
In this section, we provide additional benchmarks on the Mamba-2-Hybrid model. We begin by expanding upon Figure 2(a) with additional RULER metrics in Figure 8. Here, we observe that Mamba-2-Hybrid fine-tuned with SE-Attn consistently outperforms SW-Attn and $S^{2}$-Attn on all RULER tasks across a broad range of context sizes—even beyond the 8192 fine-tuning context size.
@@
Figure 8: Mamba-2-Hybrid RULER benchmark fine-tuned on natural language. We fine-tune Mamba-2-Hybrid with a context size of 8192 using multiple Attention variants and evaluate on the RULER benchmark (please see Sec. 11.4 for definitions of each metric). On average, fine-tuning with SE-Attn yields a performance close to the paragon model fine-tuned with Full-Attn . We observe the biggest gain in needle-in-a-haystack (NIAH) tasks, as well as variable tracking (VT), all of which require strong recall capabilities enabled by our SE-Attn.
7.1 Fine-Tuning on Natural Language + Code

### v2

@@ Appendix A Evaluating with Full Attention
Our models are efficiently fine-tuned using SE-Attn and use Full-Attn during evaluation, as in [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia]. Due to KV-caching, the complexity of Full-Attn scales linearly during evaluation, so an efficient Attention layer in this setting is not as crucial as in training. Nevertheless, for the sake of completeness, we begin by evaluating our models with the same Attention layer used during fine-tuning (i.e., we fine-tune and deploy our Hybrid model with SE-Attn). In Table 2, we experiment with models that use Full-Attn, SW-Attn, $S^{2}$-Attn, and SE-Attn; then, we evaluate perplexity on the PG-19 dataset. All models were fine-tuned with a context size of 8192. We observe that SW-Attn is best at preserving the perplexity on context sizes up to 32×32$\times$32 × larger than the one used for pre-training. SE-Attn is also able to maintain a lower perplexity across longer context sizes, while $S^{2}$-Attn deteriorates much more quickly. As discussed further in the next section, perplexity may not be the most faithful metric for assessing models’ long-context capabilities.
@@
Table 2: Mamba-2-Hybrid fine-tuned with SE-Attn or SW-Attn preserves perplexity up to 32×32$\times$32 × the pre-training context size. We fine-tune a Mamba-2-Hyrbid model (pre-trained on a context size of 2048) on a context size of 8192 using various Attention variants. We test on longer context sizes using the same Attention used during adaptation.
Appendix B LoRA for Hybrid Models and the Pitfalls of Perplexity
@@
In this section, we conduct a more thorough ablation study on the effect of training Mamba-2-Hybrid with different variants of LoRA, and provide further evidence that perplexity is not a reliable indictor of long context performance.
@@
Fine-tuning with Full-Attn. We first expand upon LABEL:fig:mamba_ruler_avg_hylora(b), where we showed that fine-tuning Mamba-2-Hybrid with SE-Attn using HyLoRA produced a stronger model than training with LoRA and LoRA+. In Figure 3, we provide a similar plot, but using Full-Attn during fine-tuning rather than SE-Attn. We observe a similar trend, where HyLoRA produces a stronger model than LoRA and LoRA+. In the remaining analyses in this section, we provide results for fine-tuning with SE-Attn, as we observed similar trends when fine-tuning with Full-Attn.
@@
Effect of LoRA rank. We next explore how the rank used for LoRA fine-tuning affects downstream performance on the RULER task. Given its stronger performance, we focus on our HyLoRA. In our previous Mamba-2-Hybrid experiments, we used a LoRA rank $(r)$ of 32 and an $\alpha$ of 64 (for our Transformer experiments, we used $r=8$ and $\alpha=16$ as in [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia]). In Figure 4, we plot the average RULER results for Mamba-2-Hybrid models fine-tuned with HyLoRA using different ranks (we scale $\alpha$ to maintain the ratio $\alpha/r=2$). Here, we observe that training with a larger rank improves downstream performance, and we start to see some saturation around $r=64$.
In Table 3, we provide PG-19 perplexity results for Mamba-2-Hybrid trained with different LoRA variants. All models are fine-tuned with a context size of 8192 and evaluated with multiple context sizes. For a given context size, we do not see a substantial difference in the perplexity. This is similar to the observation in [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia]. However, as discussed above, the rank can have a significant effect on downstream long-context tasks that require strong recall capabilities, as in RULER. Hence, while perplexity results may be promising, they are not necessarily indicative of performance on more complex long-context tasks.
@@
Figure 4: Fine-tuning with a larger LoRA rank using HyLoRA improves performance on the RULER benchmark. We fine-tune Mamba-2-Hybrid using HyLoRA with different LoRA ranks (we maintain a LoRA rank to alpha ratio of 2). We observe that fine-tuning with a larger rank produces stronger downstream results on RULER, with some saturation when using a rank of 64.
@@
Table 3: Mamba-2-Hybrid LoRA Ablations. We fine-tune Mamba-2-Hybrid with a context size of 8192 with SE-Attn using different LoRA variants. We consider LoRA, LoRA+, and HyLoRA (ours) and evaluate perplexity on the PG-19 dataset. We observe that all LoRA variants yield similar perplexity results. However, as depicted in LABEL:fig:mamba_ruler_avg_hylora and Figure 4, different LoRA variants yield substantially different performances on more complex tasks.
Appendix C Applying SE-Attention to Transformers
In this section, we show that SE-Attn can also be applied to Transformer models and use Llama 1 7B [Touvron et al.(2023a)Touvron, Lavril, Izacard, Martinet, Lachaux, Lacroix, Rozière, Goyal, Hambro, Azhar, et al.] as our representative Transformer model. When fine-tuning Llama 7B models, we found that using a fixed chunk size of $M=4096$ gave better downstream performance. All of our Llama models are fine-tuned with a context size of 16384. For SE-Attn, we use a block size of 32, and a top-$k$ of 8. For SW-Attn, we use a window size of 4096, and $S^{2}$-Attn uses the default parameters in [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia]. Due to computational constraints, we do not consider fine-tuning with Full-Attn. All Attention variants use the same RoPE [Su et al.(2024)Su, Ahmed, Lu, Pan, Bo, and Liu] scaling as in [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia].
Perplexity. Fine-tuning Llama with SE-Attn leads to better generalization on context sizes smaller and larger than the one used for fine-tuning. In Table 4, we observe that fine-tuning with SE-Attn preserves the performance of the non-fine-tuned model at smaller context sizes, and offers greater generalization to larger context sizes, as measured on PG-19 validation perplexity.
LM Harness. Fine-tuning Llama with SE-Attn yields a stronger performance on long-context tasks on the LM Evaluation Harness benchmark. As shown in Table 4, fine-tuning with SW-Attn, $S^{2}$-Attn, and SE-Attn all improve upon the non-fine-tuned model on shorter context tasks and perform similarly. However, SE-Attn gives a greater performance on long context tasks.
RULER. Fine-tuning Llama with SE-Attn produces a model with stronger performance on RULER tasks than fine-tuning with SW-Attn and $S^{2}$-Attn, as illustrated in Figure 5. On average, SW-Attn and $S^{2}$-Attn perform similarly, however, fine-tuning with SE-Attn improves performance by $\sim$5%. Despite all models having a similar PPL up to a context size of 16k as shown in Table 4, we observe a substantial difference between the RULER performance of models fine-tuned with our SE-Attn and those fine-tuned with the SW-Attn and $S^{2}$-Attn, again suggesting that PPL is not the most accurate assessment of long-context performance, as discussed in Appendix B.
Figure 5: Fine-tuning with SE-Attn outperforms SW-Attn and $S^{2}$-Attn on the RULER benchmark when applied to Llama1. We fine-tune Llama1 with a context size of 16384 using various Attention variants. We average over eleven RULER tasks, as explained in Section I.4. Fine-tuning with SE-Attn consistently outperforms SW-Attn and $S^{2}$-Attn even when evaluating on context sizes beyond the fine-tuning size.
Appendix D Beyond Mamba-2-Hybrid
In this section, we apply SE-Attn to the Zamba2 1.2B model [Glorioso et al.(2024a)Glorioso, Anthony, Tokpanov, Golubeva, Shyam, Whittington, Pilault, and Millidge]. Zamba2 uses shared Attention blocks with different LoRA adapters for the shared blocks. Despite its intricate architecture, we can replace these layers with SE-Attn and fine-tune it to perform on larger context sizes. The pre-trained model was trained with a context size of 4096 and we are able to efficiently fine-tune it with a context size of 12,288. We benchmark our model on NIAH tasks from RULER and provide results in Table 5. We compare to [Yang et al.(2025)Yang, Kautz, and Hatamizadeh] and [Yang et al.(2024b)Yang, Wang, Zhang, Shen, and Kim] and demonstrate competitive performance to SOTA models.
S-NIAH-1 S-NIAH-2 S-NIAH-3
Model 1K 2K 4K 8K 1K 2K 4K 8K 1K 2K 4K 8K
DeltaNet 97.4 96.8 99.0 98.8 98.4 45.6 18.6 14.4 85.2 47.0 22.4 -
Mamba2 99.2 98.8 65.4 30.4 99.4 98.8 56.2 17.0 64.4 47.6 4.6 -
Gated DeltaNet 98.4 88.4 91.4 91.8 100.0 99.8 92.2 29.6 86.6 84.2 27.6 -
Zamba2-Hybrid 100.0 99.2 99.4 0.0 100.0 100.0 75.0 0.0 94.4 75.4 37.6 0.0
Zamba2-Hybrid + SE-Attn 100.0 100.0 100.0 41.0 100.0 100.0 100.0 42.0 100.0 100.0 94.0 34.0
Table 5: SE-Attn can efficiently extend the context length of Zamba2-Hybrid. We fine-tune Zamab2-Hybrid using SE-Attn and evaluate on RULER NIAH tasks.
Appendix E Additional Fine-Tuning Results on Mamba-2-Hybrid
@@
In this section, we provide additional benchmarks on the Mamba-2-Hybrid model. We begin by expanding upon LABEL:fig:mamba_ruler_avg_hylora(a) with additional RULER tasks in Figure 6. Here, we observe that Mamba-2-Hybrid fine-tuned with SE-Attn consistently outperforms SW-Attn and $S^{2}$-Attn on all RULER tasks across a broad range of context sizes—even beyond the 8192 fine-tuning context size.
@@
Figure 6: Mamba-2-Hybrid RULER benchmark fine-tuned on natural language. We fine-tune Mamba-2-Hybrid with a context size of 8192 using multiple Attention variants and evaluate on the RULER benchmark (see Section I.4 for definitions of each metric). On average, fine-tuning with SE-Attn yields a performance close to the paragon model fine-tuned with Full-Attn . We observe the biggest gain in needle-in-a-haystack (NIAH) tasks, as well as variable tracking (VT), all of which require strong recall capabilities enabled by our SE-Attn.
E.1 Fine-Tuning on Natural Language + Code